Use “completing the squares” to convert the quadratic function into vertex form Given the quadratic function f(x) = 2x2– 8x + 11

Answer by ikleyn(52780)

.
f(x) = 2x^2 - 8x + 11 = 2*(x^2 - 4x) + 11 = 2*(x^2 - 2*2*x + 2^2) - 2*2^2 + 11 = 


     = 2*(x-2)^2 + 3.

Done.

--------------

To see many other similar problem solved by the  "completing the square"  method and to learn the subject,  look into the lesson
    - HOW TO complete the square - Learning by examples
in this site.